# Berkeley, CA 94720, USA

### v2

@@ 2 Geometry of Spherical Orbifolds
The $(2,0)$-theory on $X_{5}$ defines several inequivalent Hilbert spaces, and as explained in Tachikawa:2013hya , in order to specify a concrete Hilbert space one has to provide (in addition to a choice of spin structure) an element ${\mathfrak{v}}$ in the cohomology group $H^{3}(X_{5},\mathbb{Z}_{2})$. Here, the cohomology coefficient ring $\mathbb{Z}_{2}$ is identified with the center of the gauge group SU$(2)$. Compactified on $S^{1}$, the theory reduces to 5d SYM on $X_{5}$ with gauge group $\text{SO}(3)\simeq\text{SU}(2)/\mathbb{Z}_{2}$, and ${\mathfrak{v}}$ encodes the phases with which to weight, in a path integral, the contributions of topologically inequivalent $\text{SO}(3)$ gauge bundles. Specifically, the obstruction to lifting an $\text{SO}(3)$ gauge bundle on $X_{5}$ to an $\text{SU}(2)$ bundle is given by a Stiefel-Whitney class ${\mathfrak{w}}\in H^{2}(X_{5},\mathbb{Z}_{2})$, which captures the ‘t Hooft magnetic flux tHooft:1977nqb . Since $X_{5}$ is assumed orientable, we have a cohomology pairing ${\left\langle{{\mathfrak{v}}},{{\mathfrak{w}}}\right\rangle}\in\mathbb{Z}_{2}$, and the partition function of the theory labeled by ${\mathfrak{v}}$ is given, in terms of the path integrals ${\mathcal{Z}}_{\mathfrak{w}}$ that are restricted to gauge bundles with Stiefel-Whitney class ${\mathfrak{w}}$, by Tachikawa:2013hya (see also Witten:1996hc ; Freed:2006yc ; Henningson:2010rc for prior discussions related to the non uniqueness of space of the $(2,0)$ theory):
@@
where ${\mathfrak{s}}_{1},{\mathfrak{s}}_{2}\in\mathbb{Z}_{2}$, and the sign of $r^{2k}$ in the leftmost condition is required for consistency with the other two relations. Thus, the pair $({\mathfrak{s}}_{1},{\mathfrak{s}}_{2})\in\mathbb{Z}_{2}^{2}$ characterizes the obstruction to lifting an $\text{SO}(3)$ bundle over $S^{3}/\text{Dic}_{k}$ to an $\text{SU}(2)$ bundle. But since (±$(\pm r,\pm s)$, with all $4$ sign combinations, describe identical $\text{SO}(3)$ bundles, we can set ${\mathfrak{s}}_{2}=0$ for odd $k$, but we must keep it for even $k$. Therefore, for even $k$, there must be an injective map $H^{2}(S^{3}/\text{Dic}_{k},\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}^{2}$ that maps ${\mathfrak{v}}$ to $({\mathfrak{s}}_{1},{\mathfrak{s}}_{2})$, and for odd $k$ we must have an injective map $H^{2}(S^{3}/\text{Dic}_{k},\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ that maps ${\mathfrak{v}}$ to ${\mathfrak{s}}_{1}$. Indeed, from Table 4 and the universal coefficient theorem we have an isomorphism
$H^{2}(S^{3}/\text{Dic}_{k},\mathbb{Z}_{2})\cong\text{Ext}_{\mathbb{Z}}^{1}(H_{1}(S^{3}/\text{Dic}_{k},\mathbb{Z}),\mathbb{Z}_{2})=\left\{\begin{array}[]{ll}\mathbb{Z}_{2}^{2}&\text{for even $k$;}\\ \mathbb{Z}_{2}&\text{for odd $k$.}\\ \end{array}\right.$
@@
If ${\mathfrak{s}}_{2}={\mathfrak{s}}_{1}=0$ we get the solutions that can be lifted to $\text{SU}(2)$ and discussed above. Now set ${\mathfrak{s}}_{2}=0$ and ${\mathfrak{s}}_{1}=1$. Then, we diagonalize $g(r)$ and from $s^{-1}rs=-r^{-1}$ we see that either the eigenvalues of $g(r)$ are ±$\pm i$, or their product is $-1$, in which case $g(r)\notin\text{SU}(2)$. Thus, up to conjugation, there is one solution [in $\text{SU}(2)/\mathbb{Z}_{2}$]:
$g(r)=\begin{pmatrix}i&0\\ 0&-i\\ \end{pmatrix},\qquad g(s)=\pm\begin{pmatrix}e^{\frac{k\pi i}{4}}&0\\ 0&e^{-\frac{k\pi i}{4}}\\ \end{pmatrix}\qquad({\mathfrak{s}}_{1}=1).$
@@
The $S^{2}$ introduced another $\mathbb{Z}_{2}$ factor, but we note that a non-trivial Stiefel-Whitney class on $S^{2}$ does not support a flat connection, so the component of ${\mathfrak{v}}$ along $S^{2}$ has no effect on the topological partition function on $S^{3}/\text{Dic}_{k}\times S^{2}$.
@@ 3 Fermionic Chern-Simons and the Ray-Singer Torsion
Our setup, as discussed in §1, is the $(2,0)$-theory compactified on one of the 5d manifolds of §2, denoted by $X_{5}$, with a (partial) topological twist that is equivalent to identifying the $\mathfrak{so}(5)_{H}$ holonomy algebra with the $\mathfrak{so}(5)_{R}$ R-symmetry algebra. After compactification of the remaining (“time”) direction on $S^{1}$, we obtain a twisted 5d SYM theory, and we now present its field contents. In our notation, $\mu=1,\dots,5$ labels local coordinates on $X_{5}$. Due to the topological twist, the scalar fields $\Phi^{I}$ (which describe the perpendicular fluctuations of the M5555-branes in the M-theory realization, with $I=1,\dots,5$ labeling R-symmetry vector representation indices before the topological twist) get replaced by a 1-form $\Phi=\Phi_{\mu}dx^{\mu}$. The spinor fields $\psi^{a}_{\alpha}$ (with $\alpha=1,\dots,4$ a spinor index and $a=1,\dots,4$ labeling a spinor representation of the R-symmetry group), which transform in the representation $(4,4)$ of $\mathfrak{so}(5)_{H}\oplus\mathfrak{so}(5)_{R}$, decompose under the following rule:
@@
The fermionic kinetic term of the resulting 5d SYM theory can be identified with a 5d “fermionic Chern Simons theory” (FCS). In the terminology of Elliott:2020ecf (which presents a classification of topological twists of super Yang Mills theories up to 10 dimensions), the twist we are utilizing is a “$\mathcal{N}=2$ topological B twist with a supercharge of rank 4”. Unlike ordinary 3d Chern Simons theories which are constructed from connection 1-forms that are Grassmann even, the 5d FCS theories are constructed from Grassmann odd (fermionic) two-forms $\Psi$ with the action
@@
We now present the Lagrangian and the topological BRST symmetry of the twisted 5d SYM with gauge group SU$(2)$, following Bak:2017jwd . Our fields are those of Table 6, with all fields transforming in the adjoint representation of SU$(2)$. Using a point-wise linear combination of the adjoint fields, we can define twisted connections as
@@
[…]
@@
where the field strength $\mathcal{F}$ is defined with respect to the new gauge field via $\mathcal{F}=d\mathcal{A}+\mathcal{A}\wedge\mathcal{A}$, and $\epsilon$ is a Grassmann odd parameter. $\mathcal{L}_{0}$ is invariant under $\delta$, up to a total derivative, and the full Lagrangian of the twisted 5d SYM theory can be obtained by adding to $\mathcal{L}_{CS}$ a $\delta$-exact term Bak:2017jwd ,
@@
On flat space, ${\mathcal{L}}$ is the same as the Lagrangian of 5d SYM after a field redefinition is applied to the twisted fields $\chi$, $\psi$, $\rho$, $\Phi$, $\mathcal{A}$, $\overline{\mathcal{A}}$. The Lagrangian ${\mathcal{L}}$ possesses a SU$(2)$ gauge symmetry associated to the gauge field ${\mathcal{A}}$, which is gauge fixed by a standard BRST gauge fixing procedure (introducing additional Faddeev-Popov ghosts) separate from the topological BRST symmetry above.
@@
where $\mathcal{F}$ is the Euclideanized SU$(2)$ field strength. As we discussed in §2, all the five-manifolds $X_{5}$ that we will consider in this paper have $H^{1}(X_{5},\mathbb{Z})=0$, and consequently the covariant constancy condition $\mathcal{D}\Phi=0$ forces $\Phi$ to vanish since there are no nontrivial globally constant 1-forms on a manifold with trivial $H^{1}(X_{5},\mathbb{Z})$. In summary, the path integral localizes onto the moduli space of flat connections ($F=0$) with $\Phi$ vanishing.
For the manifolds $X_{5}$ that we are considering, the fundamental group is non-trivial and given by $\pi_{1}(X_{5})=\Gamma$. Therefore, there exist a finite number of non-trivial flat connections, which are classified by group homomorphisms $\Gamma\rightarrow\text{SU}(2)$ [or $\Gamma\rightarrow\text{SO}(3)$], up to conjugation. One can compute the partition function by expanding around the saddle points, which correspond to these flat connections. This approach in the context of 3d Chern-Simons theory on a general manifold is developed in Witten:1988hf ; Freed:1991wd ; Rozansky:1993zx ; Adams:1995np ; Adams:1996hi ; Marino:2002fk and reviewed in Marino:2011nm . In 3d, the renormalization of loop corrections introduces a metric-dependent phase in the partition function Witten:1988hf , but this phase is absent in 5d, and the theory remains exactly topological Bak:2017jwd .
@@
As for the fluctuations around the BPS solutions, one can show that in the limit $t\rightarrow\infty$, the action has to be expanded only up to quadratic order, and the one-loop determinant reduces to a product of functional determinants of Hodge Laplacians on 00, $1$ and $2$-forms. Note that not all quadratic terms in the action are proportional to $t$, since the $\chi\wedge d\chi$ coupling in $\mathcal{L}_{CS}$ [see (16)] is independent of $t$. However, we can follow an argument presented in Bak:2017jwd to show that (16) does not pose a problem. We first rescale all field fluctuations (both fermionic and bosonic) by $\frac{1}{\sqrt{t}}$, which results in the rescaling of $\mathcal{L}_{CS}$ by $\frac{1}{t}$. This rescaling also affects the measure of the path integral, and the factor by which the measure is rescaled can be computed by zeta-function regularization. If we expand the $p$-form fields in eigenmodes of the twisted (in the background field $A$) Hodge Laplacian $\Delta^{p}_{A}$, and consequently write the path integral measure as a product over the eigenmodes while removing the zero modes from the product, then the measure gets multiplied by a product of factors of $\sqrt{t}$ for each mode, which is regularized to
@@
where the sign (±)plus-or-minus($\pm$)( ± ) is $+$ for bosons and -- for fermions, and
@@
is the Minakshisundaram–Pleijel $\zeta$-function defined as a sum of $(-s)$-powers of the nonzero eigenvalues $\lambda_{n}$ of $\Delta^{p}_{A}$. Next, we use a result rosenberg1997laplacian that relates the value at zero of this regularized $\zeta$-function to the $p^{\text{th}}$ Betti number, which computes the rank of the twisted $p$-th cohomology group (i.e., the cohomology group obtained by replacing the exterior derivative of de-Rham cohomology with the covariant derivative with respect to the flat $\mathfrak{su}(2)$ connection $A$):
@@
[…]
@@
We still need to compute $\text{vol}(\mathcal{H}_{A})$, which is the volume of the subgroup of constant gauge transformations $\mathcal{H}_{A}$ that preserve $A$. $\mathcal{H}_{A}$ is isomorphic to a subgroup $H_{A}$ of the gauge group SU$(2)$, but its volume should be calculated using a different metric from that of SU$(2)$, and as a result it depends on the volume of the manifold $\text{vol}(X_{5})$. To see this, we will study how $\text{vol}(\mathcal{H}_{A})$ varies under rescaling of $\text{vol}(X_{5})$. $\text{vol}(\mathcal{H}_{A})$ is calculated using a metric on the tangent space of $\mathcal{H}_{A}$, which is the space of covariantly constant adjoint-valued 00-forms $\varepsilon$ on $X_{5}$, whose norm is given by
@@
For the trivial flat connection, the isotropy group is SU$(2)$, as all gauge transformations preserve the connection. For non-trivial abelian flat connections, the isotropy group is the maximal abelian subgroup of SU$(2)$, which corresponds to the maximal torus U$(1)$. For non-abelian flat connections, such as those associated with the dihedral group Dick, the isotropy group reduces to the center of SU$(2)$, which is $\mathbb{Z}_{2}$.
The second type of zero modes we must address are the zero modes of physical (non-ghost) fields. We have seen near (24), from the study of the non-trivial cohomology groups of the universal cover of $X_{5}$, that only the 00-form and $2$-form fields can have zero modes. Specifically, in the case of abelian flat connections we have fermionic zero modes associated with the 0-form $\rho$ and the 2-form $\chi$. The standard approach to address such zero modes is to insert BRST-invariant operators into the path integral, which absorb the zero modes, leading to a nonzero correlation function. However, the BRST transformations of $\rho$ and $\chi$ do not allow for simple BRST invariant operators that can soak up these modes. More importantly, there is no clear physical intuition about what types of more complicated BRST-invariant observables might be constructed to achieve this. Without any operator insertions, the partition function vanishes identically for all abelian flat connections, since those have fermionic zero modes. For example, when $X_{5}=S^{3}/\mathbb{Z}_{k}\times S^{2}$, all flat connections are abelian, as discussed in §2. It is not hard to check that loop effects cannot remove these zero modes. In the case where $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$, however, non-abelian flat connections do exist. For those non-abelian flat connections, the zero modes are lifted because there are no covariantly constant 0-forms or 2-forms. For example, if a covariantly constant $\rho$ satisfying $d_{A}\rho=0$ existed, its value $\rho(p)$ at some point $p\in X_{5}$ would have to commute with the holonomy group of $A$, which is isomorphic to a Dick subgroup of SU$(2)$, and no elements of the Lie algebra $\mathfrak{su}(2)$ are invariant under the entire group. With the zero modes removed, we can proceed to compute the partition function.
@@
$Z(X_{5})=\sum_{A}\frac{1}{\text{vol}(\mathcal{H}_{A})}\frac{\det^{\prime\frac{5}{4}}\Delta_{A}^{0}\det^{\prime\frac{1}{4}}\Delta_{A}^{2}}{\det^{\prime\frac{3}{4}}\Delta_{A}^{1}},$ (35)
In our case, only non-abelian flat connections $A$ are of interest, $\dim H_{A}=0$ in (32) (since $H_{A}$ is finite), and there are no zero modes, so $\det^{\prime}$ can be replaced with $\det$. Comparing (35) and (33), we finally arrive at:
$Z(X_{5})=\sum_{A}\frac{1}{\text{vol}(\mathcal{H}_{A})}\sqrt{\tau(X_{5},A)},$ (36)
@@
$Z(S^{3}/\Gamma\times S^{2})=\sum_{A}\tau(S^{3}/\Gamma,A),$ (38)
since $\chi(S^{2})=2$ and the Euler characteristic of the odd-dimensional manifold $S^{3}/\Gamma$ vanishes. This represents a significant simplification: to study the partition function in five dimensions, we only need to compute the torsion of the three-dimensional manifold $S^{3}/\Gamma$ over all admissible choices of flat SU($2$) connections $A$. In particular, we only need to study the finite number of flat connections which are classified by group homomorphisms $\Gamma\rightarrow$ SU($2$) [or $\Gamma\rightarrow\text{SO}(3)$], up to conjugation. In the next section, we compute these torsions explicitly.

### v3

@@ 2 Geometry of Spherical Orbifolds
The $(2,0)$-theory on $X_{5}$ defines several inequivalent Hilbert spaces, and as explained in Tachikawa:2013hya , in order to specify a concrete Hilbert space one has to provide (in addition to a choice of spin structure) an element $\mathfrak{v}$} in the cohomology group $H^{3}(X_{5},\mathbb{Z}_{2})$. Here, the cohomology coefficient ring $\mathbb{Z}_{2}$ is identified with the center of the gauge group SU(2)(2). Compactified on $S^{1}$, the theory reduces to 5d SYM on $X_{5}$ with gauge group $\text{SO}(3)\simeq\text{SU}(2)/\mathbb{Z}_{2}$, and $\mathfrak{v}$} encodes the phases with which to weight, in a path integral, the contributions of topologically inequivalent $\text{SO}(3)$ gauge bundles. Specifically, the obstruction to lifting an $\text{SO}(3)$ gauge bundle on $X_{5}$ to an $\text{SU}(2)$ bundle is given by a Stiefel-Whitney class ${\mathfrak{w}}\in H^{2}(X_{5},\mathbb{Z}_{2})$, which captures the ‘t Hooft magnetic flux tHooft:1977nqb . Since $X_{5}$ is assumed orientable, we have a cohomology pairing ${\left\langle{{\mathfrak{v}}},{{\mathfrak{w}}}\right\rangle}\in\mathbb{Z}_{2}$, and the partition function of the theory labeled by $\mathfrak{v}$} is given, in terms of the path integrals ${\mathcal{Z}}_{\mathfrak{w}}$ that are restricted to gauge bundles with Stiefel-Whitney class $\mathfrak{w}$}, by Tachikawa:2013hya (see also Witten:1996hc ; Freed:2006yc ; Henningson:2010rc for prior discussions related to the non uniqueness of space of the $(2,0)$ theory):
@@
where ${\mathfrak{s}}_{1},{\mathfrak{s}}_{2}\in\mathbb{Z}_{2}$, and the sign of $r^{2k}$ in the leftmost condition is required for consistency with the other two relations. Thus, the pair $({\mathfrak{s}}_{1},{\mathfrak{s}}_{2})\in\mathbb{Z}_{2}^{2}$ characterizes the obstruction to lifting an $\text{SO}(3)$ bundle over $S^{3}/\text{Dic}_{k}$ to an $\text{SU}(2)$ bundle. But since $(\pm r,\pm s)$, with all $4$ sign combinations, describe identical $\text{SO}(3)$ bundles, we can set ${\mathfrak{s}}_{2}=0$ for odd $k$, but we must keep it for even $k$. Therefore, for even $k$, there must be an injective map $H^{2}(S^{3}/\text{Dic}_{k},\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}^{2}$ that maps $\mathfrak{v}$} to $({\mathfrak{s}}_{1},{\mathfrak{s}}_{2})$, and for odd $k$ we must have an injective map $H^{2}(S^{3}/\text{Dic}_{k},\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ that maps $\mathfrak{v}$} to ${\mathfrak{s}}_{1}$. Indeed, from Table 4 and the universal coefficient theorem we have an isomorphism
[…]
@@
If ${\mathfrak{s}}_{2}={\mathfrak{s}}_{1}=0$ we get the solutions that can be lifted to $\text{SU}(2)$ and discussed above. Now set ${\mathfrak{s}}_{2}=0$ and ${\mathfrak{s}}_{1}=1$. Then, we diagonalize $g(r)$ and from $s^{-1}rs=-r^{-1}$ we see that either the eigenvalues of $g(r)$ are $\pm i$, or their product is 1-1, in which case $g(r)\notin\text{SU}(2)$. Thus, up to conjugation, there is one solution [in $\text{SU}(2)/\mathbb{Z}_{2}$]:
$g(r)=\begin{pmatrix}i&0\\ 0&-i\\ \end{pmatrix},\qquad g(s)=\pm\begin{pmatrix}e^{\frac{k\pi i}{4}}&0\\ 0&e^{-\frac{k\pi i}{4}}\\ \end{pmatrix}\qquad({\mathfrak{s}}_{1}=1).$ (15)
@@
The $S^{2}$ introduced another $\mathbb{Z}_{2}$ factor, but we note that a nontrivial Stiefel-Whitney class on $S^{2}$ does not support a flat connection, so the component of $\mathfrak{v}$} along $S^{2}$ has no effect on the topological partition function on $S^{3}/\text{Dic}_{k}\times S^{2}$.
@@ 3 Fermionic Chern-Simons and the Ray-Singer Torsion
Our setup, as discussed in §1, is the $(2,0)$-theory compactified on one of the 5d manifolds of §2, denoted by $X_{5}$, with a (partial) topological twist that is equivalent to identifying the $\mathfrak{so}(5)_{H}$ holonomy algebra with the $\mathfrak{so}(5)_{R}$ R-symmetry algebra. After compactification of the remaining (“time”) direction on $S^{1}$, we obtain a twisted 5d SYM theory, and we now present its field contents. In our notation, $\mu=1,\dots,5$ labels local coordinates on $X_{5}$. Due to the topological twist, the scalar fields $\Phi^{I}$ (which describe the perpendicular fluctuations of the M55-branes in the M-theory realization, with $I=1,\dots,5$ labeling R-symmetry vector representation indices before the topological twist) get replaced by a 1-form $\Phi=\Phi_{\mu}dx^{\mu}$. The spinor fields $\psi^{a}_{\alpha}$ (with $\alpha=1,\dots,4$ a spinor index and $a=1,\dots,4$ labeling a spinor representation of the R-symmetry group), which transform in the representation $(4,4)$ of $\mathfrak{so}(5)_{H}\oplus\mathfrak{so}(5)_{R}$, decompose under the following rule:
@@
The fermionic kinetic term of the resulting 5d SYM theory can be identified with a 5d “fermionic Chern Simons theory” (FCS). In the terminology of Elliott:2020ecf (which presents a classification of topological twists of super Yang Mills theories up to 10 dimensions), the twist we are utilizing is a $\mathcal{N}=2$ topological B twist with a supercharge of rank 4”. Unlike ordinary 3d Chern Simons theories which are constructed from connection 1-forms that are Grassmann even, the 5d FCS theories are constructed from Grassmann odd (fermionic) two-forms $\Psi$ with the action
@@
We now present the Lagrangian and the topological BRST symmetry of the twisted 5d SYM with gauge group SU(2)(2), following Bak:2017jwd . Our fields are those of Table 6, with all fields transforming in the adjoint representation of SU(2)(2). Using a point-wise linear combination of the adjoint fields, we can define twisted connections as
@@
$\mathcal{S}_{CS}=-\frac{i}{2}\int_{X_{5}}\operatorname{tr}\chi\wedge\mathcal{D}\chi=-\frac{i}{8}\int_{X_{5}}\epsilon^{\mu\nu\rho\sigma\tau}\operatorname{tr}\chi_{\mu\nu}\mathcal{D}_{\rho}\chi_{\sigma\tau}\hskip 2.84544ptd^{5}x.$ (20)
@@
where the field strength $\mathcal{F}$ is defined with respect to the new gauge field via $\mathcal{F}=d\mathcal{A}-i\mathcal{A}\wedge\mathcal{A}$, and $\epsilon$ is a Grassmann odd parameter. $\mathcal{L}_{0}$ is invariant under $\delta$, up to a total derivative, and the full Lagrangian of the twisted 5d SYM theory can be obtained by adding to $\mathcal{L}_{CS}$ a $\delta$-exact term Bak:2017jwd ,
@@
On flat space, $\mathcal{L}$} is the same as the Lagrangian of 5d SYM after a field redefinition is applied to the twisted fields $\chi$, $\psi$, $\rho$, $\Phi$, $\mathcal{A}$, $\overline{\mathcal{A}}$. The Lagrangian $\mathcal{L}$} possesses a SU(2)(2) gauge symmetry associated to the gauge field $\mathcal{A}$}, which is gauge fixed by a standard BRST gauge fixing procedure (introducing additional Faddeev-Popov ghosts) separate from the topological BRST symmetry above.
@@
where $\mathcal{F}$ is the Euclideanized SU(2)(2) field strength. As we discussed in §2, all the five-manifolds $X_{5}$ that we will consider in this paper have $H^{1}(X_{5},\mathbb{Z})=0$, and consequently the covariant constancy condition $\mathcal{D}\Phi=0$ forces $\Phi$ to vanish since there are no nontrivial globally constant 1-forms on a manifold with trivial $H^{1}(X_{5},\mathbb{Z})$. In summary, the path integral localizes onto the moduli space of flat connections ($F=0$) with $\Phi$ vanishing.
For the manifolds $X_{5}$ that we are considering, the fundamental group is nontrivial and given by $\pi_{1}(X_{5})=\Gamma$. Therefore, there exist a finite number of nontrivial flat connections, which are classified by group homomorphisms $\Gamma\rightarrow\text{SU}(2)$ [or $\Gamma\rightarrow\text{SO}(3)$], up to conjugation. One can compute the partition function by expanding around the saddle points, which correspond to these flat connections. This approach in the context of 3d Chern-Simons theory on a general manifold is developed in Witten:1988hf ; Freed:1991wd ; Rozansky:1993zx ; Adams:1995np ; Adams:1996hi ; Marino:2002fk and reviewed in Marino:2011nm . In 3d, the renormalization of loop corrections introduces a metric-dependent phase in the partition function Witten:1988hf , but this phase is absent in 5d, and the theory remains exactly topological Bak:2017jwd .
@@
As for the fluctuations around the BPS solutions, one can show that in the limit $t\rightarrow\infty$, the action has to be expanded only up to quadratic order, and the one-loop determinant reduces to a product of functional determinants of Hodge Laplacians on 0, 11 and $2$-forms. Note that not all quadratic terms in the action are proportional to $t$, since the $\chi\wedge d\chi$ coupling in $\mathcal{L}_{CS}$ [see (20)] is independent of $t$. However, we can follow an argument presented in Bak:2017jwd to show that (20) does not pose a problem. We first rescale all field fluctuations (both fermionic and bosonic) by $\frac{1}{\sqrt{t}}$, which results in the rescaling of $\mathcal{L}_{CS}$ by $\frac{1}{t}$. This rescaling also affects the measure of the path integral, and the factor by which the measure is rescaled can be computed by zeta-function regularization. If we expand the $p$-form fields in eigenmodes of the twisted (in the background field $A$) Hodge Laplacian $\Delta^{p}_{A}$, and consequently write the path integral measure as a product over the eigenmodes while removing the zero modes from the product, then the measure gets multiplied by a product of factors of $\sqrt{t}$ for each mode, which is regularized to
@@
where the sign $(\pm)$ is ++ for bosons and - for fermions, and
@@
is the Minakshisundaram–Pleijel $\zeta$-function defined as a sum of (s)(-s)-powers of the nonzero eigenvalues $\lambda_{n}$ of $\Delta^{p}_{A}$. Next, we use a result rosenberg1997laplacian that relates the value at zero of this regularized $\zeta$-function to the $p^{\text{th}}$ Betti number, which computes the rank of the twisted $p$-th cohomology group (i.e., the cohomology group obtained by replacing the exterior derivative of de-Rham cohomology with the covariant derivative with respect to the flat $\mathfrak{su}(2)$ connection $A$):
@@
Our theory has fields that are 0-, 1-, and 2-forms, so we are only interested in $p=0,1,2$, and for the manifolds $X_{5}$ that we consider $b_{1}(A)=0$. [^3] This statement is independent of the flat connection $A$, because with a gauge transformation we can convert a zero mode of the twisted Hodge Laplacian on $X_{5}$ to a zero mode of the standard Laplacian (with $A=0$) on a simply connected cover of $X_{5}$, and all the manifolds that we consider have finite simply connected covers. Thus, the total measure gets multiplied by
@@
We still need to compute $\text{vol}(\mathcal{H}_{A})$, which is the volume of the subgroup of constant gauge transformations $\mathcal{H}_{A}$ that preserve $A$. $\mathcal{H}_{A}$ is isomorphic to a subgroup $H_{A}$ of the gauge group SU(2)(2), but its volume should be calculated using a different metric from that of SU(2)(2), and as a result it depends on the volume of the manifold $\text{vol}(X_{5})$. To see this, we will study how $\text{vol}(\mathcal{H}_{A})$ varies under rescaling of $\text{vol}(X_{5})$. $\text{vol}(\mathcal{H}_{A})$ is calculated using a metric on the tangent space of $\mathcal{H}_{A}$, which is the space of covariantly constant adjoint-valued 0-forms $\varepsilon$ on $X_{5}$, whose norm is given by
@@
For the trivial flat connection, the isotropy group is SU(2)(2), as all gauge transformations preserve the connection. For nontrivial abelian flat connections, the isotropy group is the maximal abelian subgroup of SU(2)(2), which corresponds to the maximal torus U(1)(1). For non-abelian flat connections, such as those associated with the dihedral group Dick, the isotropy group reduces to the center of SU(2)(2), which is $\mathbb{Z}_{2}$.
The second type of zero modes we must address are the zero modes of physical (non-ghost) fields. We have seen near (28), from the study of the nontrivial cohomology groups of the universal cover of $X_{5}$, that only the 0-form and $2$-form fields can have zero modes. Specifically, in the case of abelian flat connections we have fermionic zero modes associated with the 0-form $\rho$ and the 2-form $\chi$. The standard approach to address such zero modes is to insert BRST-invariant operators into the path integral, which absorb the zero modes, leading to a nonzero correlation function. However, the BRST transformations of $\rho$ and $\chi$ do not allow for simple BRST invariant operators that can soak up these modes. More importantly, there is no clear physical intuition about what types of more complicated BRST-invariant observables might be constructed to achieve this. Without any operator insertions, the partition function vanishes identically for all abelian flat connections, since those have fermionic zero modes. For example, when $X_{5}=S^{3}/\mathbb{Z}_{k}\times S^{2}$, all flat connections are abelian, as discussed in §2. It is not hard to check that loop effects cannot remove these zero modes. In the case where $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$, however, non-abelian flat connections do exist. For those non-abelian flat connections, the zero modes are lifted because there are no covariantly constant 0-forms or 2-forms. For example, if a covariantly constant $\rho$ satisfying $d_{A}\rho=0$ existed, its value $\rho(p)$ at some point $p\in X_{5}$ would have to commute with the holonomy group of $A$, which is isomorphic to a Dick subgroup of SU(2)(2), and no elements of the Lie algebra $\mathfrak{su}(2)$ are invariant under the entire group. With the zero modes removed, we can proceed to compute the partition function.
@@
$Z(X_{5})={\mathcal{C}}\sum_{A}\frac{1}{\text{vol}(\mathcal{H}_{A})}\frac{\det^{\prime\frac{5}{4}}\Delta_{A}^{0}\det^{\prime\frac{1}{4}}\Delta_{A}^{2}}{\det^{\prime\frac{3}{4}}\Delta_{A}^{1}},$ (39)
where $\mathcal{C}$} is a normalization constant which usually is fixed by imposing that trivial flat connection gives partition function equal to unity. (See also Murthy:2025ioh for a recent discussion of normalization of the measure in the context of localization.) Since for us trivial flat connection does not contribute, we fix this normalization to be ${\mathcal{C}}=1$ using a different argument that will be presented in §5. In our case, only non-abelian flat connections $A$ are of interest, $\dim H_{A}=0$ in (36) (since $H_{A}$ is finite), and there are no zero modes, so $\det^{\prime}$ can be replaced with $\det$. Comparing (39) and (37), we finally arrive at:
$Z(X_{5})={\mathcal{C}}\sum_{\text{nonab $A$}}\frac{\sqrt{\tau(X_{5},A)}}{|H_{A}|},$ (40)
@@
$Z(S^{3}/\Gamma\times S^{2})={\mathcal{C}}\sum_{\text{nonab $A$}}\frac{\tau(S^{3}/\Gamma,A)}{|H_{A}|},$ (42)
since $\chi(S^{2})=2$ and the Euler characteristic of the odd-dimensional manifold $S^{3}/\Gamma$ vanishes. This represents a significant simplification: to study the partition function in five dimensions, we only need to compute the torsion of the three-dimensional manifold $S^{3}/\Gamma$ over all admissible choices of flat SU(2)(2) connections $A$. In particular, we only need to study the finite number of flat connections which are classified by group homomorphisms $\Gamma\rightarrow$ SU(2)(2) [or $\Gamma\rightarrow\text{SO}(3)$], up to conjugation. In the next section, we compute these torsions explicitly.